OpenStudy (anonymous):

I need help please.... Find the domain of the rational function F(x)=-2(x^2-4)/3(x^2+4x+4) I have factored them correctly but it said after I factored them I shouldn't reduce any common factors and should find the domain of the function using its factored form and that I need to state the domain in interval notation. I am lost and do not know how to work this out apparently

OpenStudy (nenadmatematika):

\[x^2+4x+4\neq0 \] \[(x+2)^2\neq0\] so the domain interval is:

OpenStudy (nenadmatematika):

\[x^2+4x+4\neq0 \] \[(x+2)^2\neq0\] so the domain interval is: \[x \in(-\infty,-2)\cup(-2,+\infty)\]

OpenStudy (anonymous):

R-{-2}

OpenStudy (anonymous):

Check your nos nena..

OpenStudy (nenadmatematika):

to check what?

OpenStudy (anonymous):

I am not understanding...i was told to find the domain of the function using the factored form F(x)=[-2(x-2)(x+2)]/[3(x+2)(x+2)]....how do I work this out

OpenStudy (anonymous):

Sorry my bad

OpenStudy (anonymous):

Once factored you must use fact the Denominator cannot be zero for any real no

OpenStudy (nenadmatematika):

it's ok...:D kcbrosell what part you don't understand?

OpenStudy (anonymous):

all of it I am failing this class big time...I dont understand all the factoring or how to put it into and equation and so on

OpenStudy (nenadmatematika):

every time you want to find the domain of the function which is given in fraction form denominator must not be equal to zero, so when you factor x^2+4x+4 you get the result I wrote.....

OpenStudy (anonymous):

I am not understanding how you get to the result u wrote

OpenStudy (nenadmatematika):

well it's a basic binomial formula which you must know if you want to do some serious math in the future....

OpenStudy (anonymous):

I got to this point x^2-4=(x-2)(x+2) & x^2+4x+4=(x^2+2)

OpenStudy (nenadmatematika):

you got wrong x^2+2x+4...it is (x+2)^2

OpenStudy (anonymous):

o crud...so after I get that worked out do I have to put them into another equation

OpenStudy (anonymous):